Q.1 Consider t

(d) Determine the equilibrium levels of all endogenous variables under the assumptions of the classical

macroeconomic framework.

(e) Beginning from the initial classical equilibrium, suppose that the central bank increases the money supply

by 420 while price remains fixed at its initial long run equilibrium level. What will be the impact of this

policy on all endogenous variables in short run and long run?

(i) Starting from the initial equilibrium position again, suppose that the capital stock increases by 170. What

will be the impact of the expansion on labour market equilibrium and aggregate supply of output?

Calculate values of all endogenous variables and give intuitive explanation of the results.

(j) Compare the equilibrium positions in (d) and (i) indicating all points.

(k) Suppose that Li → ∞ in Equation.9 of the model. How will it affect the shape of the money demand and

the LM curve. Will the monetary policy of part (e) have the same


1
Expert's answer
2021-07-20T10:45:08-0400

The complete question is:

Q.1 Consider the following information about a hypothetical economy:

1. Y = A ( ) 0.025K − 0.5N N

2. A=2/3

3. K = 2000

4. N^s=-18+(18/5)w

5. C=200+(2/3)(Y-T)-300r

6. T=-75+(1/4)Y

7. I =100−100r

8. G =100

9. L = 0.5Y − 200i

10. M = 6300

11. 0.10


d).


"Y=\frac{2}{3}[(0.025×2000)-0.5N]N\\Y=\frac{2}{3}(50-0.5N)N\\Y=\frac{100N}{3}-\frac{1}{3}N^2"

equilibrium income


"T=-75+\frac{1}{4}Y\\T=-75+\frac{25y}{3}-\frac{y^3}{12}"

taxes


"C=300+\frac{2}{3}(Y-T)-300r\\C=300+\frac{2}{3}(Y-(-75+\frac{1}{4}y))-300r\\C=-300r+\frac{y}{2}+250\\C=-300r+250+\frac{\frac{100N}{3}-\frac{N^3}{3}}{2}"

consumption


"N=-18+\frac{18}{5}W\\Y=\frac{594}{5}W-594\\Y-594=\frac{594}{5}W\\W=\frac{5(Y-594)}{594}\\N=-18+\frac{18}{5}(\frac{5(Y-594)}{594})\\N=\frac{Y-1188}{33}"


"I=100-100r\\r=\frac{1000-r}{100}\\interest \space rates\\"


any change in in values affects each other because they are correlated. the economy depends on several factors to be at equilibriu


e)In the short run since the price remain fixed, the central bank might simultaneously engage in expansionary monetary policy to lower the interest rate to the initial rates while in the long-run price will fall over time, and the economy returns to its normal interest rates


i)The Capital stock increase will automatically raises the labour market equilibrium as wages will increase. At this point, labour market tend to adjust towards an equilibrium having higher wages and employment amount. Further, the increase in capital stock contributes to higher potential output level, hence aggregate supply of output shifting towards right.


j)g) Expansion policy will lead to increased money supply through increased government spending and tax cuts. This is because these factors influence employment and household income.

h)




k)Expansion policy will lead to increased money supply through increased government spending and tax cuts. This is because these factors influence employment and household income.